Solve the system by addition/subtraction:
-x = -2y - 16
x + y = 4

Answer by amarjeeth123(569)

The given equations are
-x = -2y - 16..........equation 1
x + y = 4..............equation 2
Adding 1 and 2 we get,
-x+x+y=-2y-16+4
y+2y=-12
3y=-12
y=-4
x=8